OpenStudy (anonymous):

Solve the system using any algebraic method. (I used matrices and I got 0,0,1??!! Medal & fan for help!! x+2y+3z=5 3y-4z=-9 x=-2z

OpenStudy (annonomys):

Is that the Equation?

OpenStudy (anonymous):

yes

OpenStudy (annonomys):

Mk.

OpenStudy (annonomys):

Let's solve for x. x+2y+3z=5 Step 1: Add -2y to both sides. x+2y+3z+−2y=5+−2y x+3z=−2y+5 Step 2: Add -3z to both sides. x+3z+−3z=−2y+5+−3z x=−2y−3z+5

OpenStudy (anonymous):

ok

OpenStudy (annonomys):

Did that help?

OpenStudy (anonymous):

Let me try solving it first...

OpenStudy (anonymous):

I got 5, -1.59, 1.06. It looks about right I rounded it to the nearest hundreths

OpenStudy (johnweldon1993):

\[\large x + 2y + 3z = 5\] \[\large 3y - 4z = -9\] \[\large x = -2z\] We can solve that second equation for 'y' and substitute it into the first equation \[\large 3y - 4z = -9 \rightarrow y = \frac{-9 + 4z}{3}\] Replace that in the first as well as replace 'x' with -2z *from the third equation* \[\large x + 2y + 3z = 5 \rightarrow -2z + 2(\frac{-9 + 4z}{3}) + 3z = 5\] \[\large -2z + \frac{-18 + 8z}{3} + 3z = 5\] \[\large -6z -18 + 8z + 9z = 15\] \[\large 11z = 33\] \[\large z = \frac{33}{11} = 3\] From there...you can solve for 'x'...and then for 'y'

OpenStudy (anonymous):

Thank you sir!

OpenStudy (johnweldon1993):

No problem, but let me know if anything doesnt make sense!

OpenStudy (anonymous):

Wait, I try pluggin in the numbers but it doesn't match up

OpenStudy (anonymous):

Do you mind pluggin it in for me pls?

OpenStudy (johnweldon1993):

Well what do you get for the numbers? z = 3...what do you get for x and y?

OpenStudy (anonymous):

x=-6 and y=5?

OpenStudy (johnweldon1993):

My 'x' matches but not 'y' Lets see...

OpenStudy (anonymous):

ok

OpenStudy (annonomys):

Good Job, Jasmine.

OpenStudy (anonymous):

Thank you!

OpenStudy (johnweldon1993):

So just take any of the equations...like the first one \[\large x + 2y + 3z = 5\] \[\large -6 + 2y + 3(3) = 5\] \[\large -6 + 2y + 9 = 5\] \[\large 3 + 2y = 5\] \[\large 2y = 5 - 3\] \[\large 2y = 2\] \[\large y = 1\]

OpenStudy (annonomys):

You Are Welcome!

OpenStudy (anonymous):

Thx @johnweldon1993. I see where I messed up

OpenStudy (johnweldon1993):

Not a problem :)

OpenStudy (anonymous):

Can u pls help me with the others too?

OpenStudy (anonymous):

I can't close my current problem so i am just going to post it here. -4x-y+5z=-7 2y+18z=0 2x+y+2z=5

OpenStudy (annonomys):

Mk. i will help you.

OpenStudy (anonymous):

Thanks :)

OpenStudy (annonomys):

No Problem.

OpenStudy (anonymous):

@braydenbunner

OpenStudy (anonymous):

Help pls

OpenStudy (johnweldon1993):

Oh sorry didnt realize you posted back lol

OpenStudy (anonymous):

it's totally fine!

OpenStudy (annonomys):

So, Whats your next question?

OpenStudy (anonymous):

I can't close my current problem so i am just going to post it here. -4x-y+5z=-7 2y+18z=0 2x+y+2z=5

OpenStudy (annonomys):

That is ok.

OpenStudy (annonomys):

Ok for the first one. What are you solving for X or Y?

OpenStudy (anonymous):

i think y?

OpenStudy (annonomys):

Alright Well. Let's solve for y. −4x−y+5z=−7 Step 1: Add 4x to both sides. −4x−y+5z+4x=−7+4x −y+5z=4x−7 Step 2: Add -5z to both sides. −y+5z+−5z=4x−7+−5z −y=4x−5z−7 Step 3: Divide both sides by -1. −y −1 = 4x−5z−7 −1 y=−4x+5z+7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can i see it with x

OpenStudy (johnweldon1993):

You'll find the answer to this one quick :)

OpenStudy (annonomys):

Ok.

OpenStudy (anonymous):

ok, but i am struggling..

OpenStudy (johnweldon1993):

\[\large -4x-y+5z=-7\] \[\large 2y+18z=0\] \[\large 2x+y+2z=5\] Lets add the first and third equations..it will eliminate the 'y' \[\large -4x - y + 5z = -7\] \[\large +(2x + y + 2x = 5)\] Gives us \[\large -2x + 7z = -2\] Now we only have 2 equations to work with... \[\large -2x + 7z = -2\] and \[\large 2y + 18z = 0\] What do you notice about these 2 equations?

OpenStudy (annonomys):

Let's solve for x. −4x−y+5z=−7 Step 1: Add y to both sides. −4x−y+5z+y=−7+y −4x+5z=y−7 Step 2: Add -5z to both sides. −4x+5z+−5z=y−7+−5z −4x=y−5z−7 Step 3: Divide both sides by -4. −4x −4 = y−5z−7 −4 x= −1 4 y+ 5 4 z+ 7 4

OpenStudy (anonymous):

one of them is lacking a x and y. also the both have a z

OpenStudy (johnweldon1993):

Well even more general....in total with 2 equations...we have 3 different unknowns we have to solve for And how could we POSSIBLY solve 3 unknowns with only 2 equations...we cannot This system of equations has no solution

OpenStudy (anonymous):

oh ok. no wonder why i couldn't solve it

OpenStudy (anonymous):

so would i just put no solution?

OpenStudy (johnweldon1993):

Yeah, no solution :)

OpenStudy (anonymous):

thanks @johnweldon1993

OpenStudy (annonomys):

Thank you johnweldon1993. Much Thanks

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

omg my computer is being slow

OpenStudy (anonymous):

I have more that I need help with -x-y-2z=9 -2x-2y-z=-1 -z-y+z=-10

OpenStudy (anonymous):

ooops i meant 1 not =-1

OpenStudy (anonymous):

-x-y-2z=9 -2x-2y-z=1 -z-y+z=-10

OpenStudy (johnweldon1993):

Is that third equation correct?

OpenStudy (anonymous):

oops it's wrong. it is suppose to be -x

OpenStudy (johnweldon1993):

Damn I was gonna say this got a whole lot easier lol

OpenStudy (anonymous):

lol

OpenStudy (annonomys):

It gets easier as you go along and learn the steps.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@johnweldon1993 pls help

OpenStudy (annonomys):

What do you need help with?

OpenStudy (anonymous):

with that. I need explanations.

OpenStudy (annonomys):

Mk.

OpenStudy (annonomys):

Let's solve for x. −2x−2y−z=1 Step 1: Add 2y to both sides. −2x−2y−z+2y=1+2y −2x−z=2y+1 Step 2: Add z to both sides. −2x−z+z=2y+1+z −2x=2y+z+1 Step 3: Divide both sides by -2. −2x −2 = 2y+z+1 −2 x=−y+ −1 2 z+ −1 2

OpenStudy (annonomys):

i typd it up while i was sitting here.

OpenStudy (anonymous):

ok so x=2

OpenStudy (annonomys):

Yep Good Job! :)

OpenStudy (anonymous):

thanks! can you help show me how you solve for y and z too pls?

OpenStudy (annonomys):

Sure.

OpenStudy (annonomys):

Let's solve for z. −2x−2y−z=1 Step 1: Add 2x to both sides. −2x−2y−z+2x=1+2x −2y−z=2x+1 Step 2: Add 2y to both sides. −2y−z+2y=2x+1+2y −z=2x+2y+1 Step 3: Divide both sides by -1. −z −1 = 2x+2y+1 −1 z=−2x−2y−1

OpenStudy (anonymous):

Thanks!

OpenStudy (johnweldon1993):

Not to interrupt *And sorry I was away* I'm not getting a solution here...I'm gonna check it again but

OpenStudy (anonymous):

It's totally fine. Ok

OpenStudy (annonomys):

Let's solve for y. −2x−2y−z=1 Step 1: Add 2x to both sides. −2x−2y−z+2x=1+2x −2y−z=2x+1 Step 2: Add z to both sides. −2y−z+z=2x+1+z −2y=2x+z+1 Step 3: Divide both sides by -2. −2y −2 = 2x+z+1 −2 y=−x+ −1 2 z+ −1 2

OpenStudy (anonymous):

so y=2?

OpenStudy (annonomys):

Yes Good Job!! :)

OpenStudy (anonymous):

Thx. can u pls show me how to solve for z

OpenStudy (annonomys):

Sure thing.

OpenStudy (annonomys):

Let's solve for z. −2x−2y−z=1 Step 1: Add 2x to both sides. −2x−2y−z+2x=1+2x −2y−z=2x+1 Step 2: Add 2y to both sides. −2y−z+2y=2x+1+2y −z=2x+2y+1 Step 3: Divide both sides by -1. −z −1 = 2x+2y+1 −1 z=−2x−2y−1

OpenStudy (anonymous):

The answers don't match up quite well

OpenStudy (annonomys):

I know.

OpenStudy (anonymous):

@Directrix pls help me

OpenStudy (annonomys):

Its the number that is the answer. If you get it wrong let me know and i will Explain it in another way.

OpenStudy (anonymous):

here can you try plugin in the variables.

OpenStudy (anonymous):

i want to see if i mess up somewhere

OpenStudy (annonomys):

Mk

OpenStudy (anonymous):

@pooja195 pls help, I am struggling.

OpenStudy (annonomys):

@pooja195 Ma'am Is it ok if you help this young lady?

OpenStudy (anonymous):

@satellite73 pls help

OpenStudy (annonomys):

It seams they are offline :/.

OpenStudy (anonymous):

yea

OpenStudy (annonomys):

Do you have anymore?

OpenStudy (anonymous):

yea